OpenStudy (anonymous):

P(x)=x3−5x2−4x+20 factor this polynomial

OpenStudy (asnaseer):

\[P(x) = x ^{3} - 5x ^{2} - 4x + 20 = x ^{2} (x - 5) - 4x + 20 = x ^{2}(x-5) - 4(x - 5)\] \[P(x) = (x ^{2} - 4)(x-5) = (x-2)(x+2)(x-5)\]

hero (hero):

If someone presents a solution to your problem, and you don't have any other questions, feel free to give a medal